Please complete full question (A & B) for immediate upvote.

A) Review the characteristics of sinusoidal functions. As you do so, remember that the goal of this course is to relate mathematical models—such as sinusoidal functions—to physical circuits and the signals that flow through them. So, what are the four quantitative characteristics of sinusoids that we need to know to completely specify a sinusoidal function mathematically? I.e. if we want to express an AC voltage, v(t), in a circuit as a sinusoid, what quantities do we need to know in order to write v(t) = f[sin(t)], where f represents a particular expression of a sine function?

B) Next, imagine that you’ve got a graph of that same v(t) sitting on the lab bench in front of you. How are you going to figure out, using the graph, all of the quantitative information you need to write down the mathematical function f? In particular, what are the things you need to know about the graph in order to find f?
Answer #1

A.) Quantitative characteristics of sinusoids  are 1. maximum amplitude and 2. time period or frequency or angular frequency

T= 1/f = 2/w 3. phase angle

B.) first of all we have to see the maximum amplitude of sinusoid and then time period - time interval at which same graph continues. then we can determine the 'f'. generally sine function starts at origin with '0' value but many functions don't because they have phase angle

as the above matter may confuse you so i will give information with graph. ( i had got this graph google, so i pasted here, if you want indepth details, comment me)